OpenStudy (anonymous):

How do you convert 3 moles to molcules?

OpenStudy (abhisar):

Hi! 1 mole of molecules means \(\sf 6.023 \times 10^{23}\) molescules. So 3 moles of molecules will equal to \(\sf 3 \times 6.023 \times 10^{23}\) molecules.

OpenStudy (anonymous):

Just Use avogadros # which is 6.023×1023 for yor question it would be 3x6.023×1023

OpenStudy (arindameducationusc):

I agree with @abhisar.... 1 mole of molecules means 6.023×10^23 molescules. So 3 moles of molecules will equal to 3×6.023×10^23 molecules. Its like 12 bananas= 1 dozen 6.022*10^23 molecules or bananas= 1 mole (Its just a number ) :D